OpenStudy (anonymous):

PROVE: 1/secx-1 + 1/secx+1 = 2/cos tan^2x

OpenStudy (campbell_st):

well you need a common denominator... so cross multiply the for the numerators and get a common denominator \[\frac{1}{\sec(x) -1} + \frac{1}{\sec(x) +1} = \frac{\sec(x) + 1 + \sec(x) - 1}{(\sec(x) -1)(\sec(x) + 1)}\] now the numerator simplifies and the denominator is the difference of 2 squares \[\frac{2 \sec(x)}{\sec^2(x) - 1}\] now all yo need to do is realise sec = 1/cos and sec^2 - 1 = tan^2 substitute them and you'll find your answer

OpenStudy (anonymous):

You're supposed to solve for one of them only to get to the other....so I don't understand what you just did...

OpenStudy (campbell_st):

well I simplified the left hand side, basically I made it possible to add the 2 fractions... now you need to make some substitutions to the left hand side

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

i get it...thanks